Ferris Company began January with 6,000 units of its principal product. The cost of each unit is $6. Merchandise transactions for the month of January are as follows:

Purchases Date of Purchase Units Units Unit Cost* Total Cost

Jan. 10 5,000 $ 7 $ 35,000

Jan. 18 6,000 $8 $48,000

11,000 $83,000

* Includes purchase price and cost of freight. Sales Date of Sale Units Jan. 5 3,000 Jan. 12 2,000 Jan. 20 4,000 Total 9,000 8,000 units were on hand at the end of the month.

1. Calculate January's ending inventory and cost of goods sold for the month using FIFO, periodic system.

2. Calculate January's ending inventory and cost of goods sold for the month using LIFO, periodic system.

3. Calculate January's ending inventory and cost of goods sold for the month using FIFO, perpetual system.
Answer #1

1) Calculate following

Ending inventory under FIFO = 6000*8+2000*7 = 62000

Cost of goods sold under FIFO = 119000-62000 = 57000

2) Calculate following

Ending inventory under LIFO = 6000*6+2000*7 = 50000

Cost of goods sold under LIFO = 119000-50000 = 69000

3) Calculate following

Ending inventory under FIFO = 6000*8+2000*7 = 62000

Cost of goods sold under FIFO = 119000-62000 = 57000